# AgentHER: Hindsight Experience Replay for LLM Agent Trajectory Relabeling

### 22 Mar 2026 v1

@@ 1 Introduction
  • Novel formulation: first systematic pipeline bridging goal-conditioned HER to natural-language agent data synthesis (as opposed to inference-time memory, e.g. ECHO; Hu et al., 2025), with explicit treatment of goal-text faithfulness and factual grounding that distinguishes it from prior failure-reuse and goal-relabeling work (see Section 2).
@@ 2 Related Work
HER (Andrychowicz et al., 2017) converts sparse-reward RL failures by substituting the achieved state as the new goal (Kaelbling, 1993; Sutton et al., 1999; Mnih et al., 2015). Lifting HER to LLM agents requires understanding what was achieved across unstructured multi-step tool interactions, and synthesising a natural-language prompt that the trajectory genuinely satisfies—two capabilities absent from prior RL formulations. Existing failure-reuse work in NLP selects or reweights (goal, trajectory) pairs (Peng et al., 2023; Gulcehre et al., 2023; Hosseini et al., 2024) but never generates a new achievable goal; that reverse-engineering step is AgentHER’s core novelty. Concurrently, Hu et al. (2025) propose ECHO (Experience Consolidation via Hindsight Optimization), a prompting framework that adapts HER for online LM agents: the model identifies alternative subgoals from failed trajectories and maintains optimized trajectory descriptions in a scratchpad memory to improve sample efficiency at inference time. AgentHER differs in targeting offline training data augmentation: we relabel only the goal (user prompt) while keeping the trajectory unchanged, and output SFT/DPO datasets for fine-tuning rather than inference-time memory; the two approaches are complementary.

### 25 Apr 2026 v3

@@ 1 Introduction
  • Novel formulation: a systematic pipeline bridging goal-conditioned HER to natural-language agent data synthesis (as opposed to inference-time memory, e.g. ECHO; Hu et al., 2025), with explicit treatment of goal-text faithfulness and factual grounding that distinguishes it from concurrent hindsight relabeling work (HSL; Li et al., 2026) and prior failure-reuse and goal-relabeling work (see Section 2). Unlike HSL, AgentHER targets exclusively offline-collected failures and relabels the top-level goal with multi-judge verification.
@@ 2 Related Work
HER (Andrychowicz et al., 2017) converts sparse-reward RL failures by substituting the achieved state as the new goal (Kaelbling, 1993; Sutton et al., 1999; Mnih et al., 2015). Lifting HER to LLM agents requires understanding what was achieved across unstructured multi-step tool interactions, and synthesising a natural-language prompt that the trajectory genuinely satisfies—two capabilities absent from prior RL formulations. Existing failure-reuse work in NLP selects or reweights (goal, trajectory) pairs (Peng et al., 2023; Gulcehre et al., 2023; Hosseini et al., 2024) but never generates a new achievable goal; that reverse-engineering step is AgentHER’s core novelty. Concurrently, Hu et al. (2025) propose ECHO (Experience Consolidation via Hindsight Optimization), a prompting framework that adapts HER for online LM agents: the model identifies alternative subgoals from failed trajectories and maintains optimized trajectory descriptions in a scratchpad memory to improve sample efficiency at inference time. AgentHER differs in targeting offline training data augmentation: we relabel only the goal (user prompt) while keeping the trajectory unchanged, and output SFT/DPO datasets for fine-tuning rather than inference-time memory; the two approaches are complementary. Also concurrently, Li et al. (2026) propose Hindsight Supervised Learning (HSL), which similarly uses an auxiliary LLM to relabel agent trajectories in hindsight; HSL mines all goals achieved along a trajectory (including intermediate subgoals during successful runs) and maintains a continuously refreshed on-policy buffer, while adding irrelevant-action masking and demonstration reweighting. AgentHER is complementary: it focuses exclusively on failed trajectories from offline collections, relabels the top-level goal while preserving the full trajectory, introduces multi-judge verification and failure-severity weighting to control label noise, and targets WebArena/ToolBench web-navigation and API-use settings rather than the embodied-agent benchmarks (ALFWorld, PlanCraft, WebShop) studied in HSL.